OpenStudy (abbycross167):

Can someone please help me? I really need help... Question~ Find the point-slope form of the equation that goes through these two points. (1, 20) and (8, 4.5)

OpenStudy (bundleoftea0196):

The point - slope equation has a general from of (y-y1)=m(x-x1) where x1, y1 are chosen points belonging to the graph and m is the slope of the linear equation. Find the slope from point (1, 20) to point (8, 4.5) and you have your m. Then choose either of those points as your (x1, y1) and plug it into the general form.

OpenStudy (abbycross167):

@bundleoftea0196 like this? Y= y2-y1/x2-x1 (4.5-20)/(8-1)

OpenStudy (abbycross167):

@jim_thompson5910 do you think you could try to help me?

jimthompson5910 (jim_thompson5910):

you have the right idea on slope so far, you just need to compute `(4.5-20)/(8-1)`

OpenStudy (abbycross167):

how do I do that?

jimthompson5910 (jim_thompson5910):

what is `4.5-20` equal to?

OpenStudy (abbycross167):

15.5?

jimthompson5910 (jim_thompson5910):

-15.5 you mixed up the order

OpenStudy (abbycross167):

oh okay, then what?

jimthompson5910 (jim_thompson5910):

8-1 is equal to 7, so we have `(4.5-20)/(8-1)` turn into `-15.5/7` which is equal to what?

OpenStudy (abbycross167):

-31/2?

jimthompson5910 (jim_thompson5910):

use a calculator to compute -15.5/7 it will be some decimal number

OpenStudy (abbycross167):

-2.214?

jimthompson5910 (jim_thompson5910):

correct, that's the approximate slope

jimthompson5910 (jim_thompson5910):

now go to the equation @bundleoftea0196 wrote out

jimthompson5910 (jim_thompson5910):

Let's replace m with -2.214, since we now know the approx slope is -2.214 \[\Large y - y_1 = m(x - x_1)\] \[\Large y - y_1 = {\color{red}{m}}(x - x_1)\] \[\Large y - y_1 = {\color{red}{-2.214}}(x - x_1)\]

jimthompson5910 (jim_thompson5910):

The \(\Large x_1\) and \(\Large y_1\) will be replaced with the x coordinate and y coordinate (respectively) of exactly one of the points this line goes through. So you just pick one of the points, either (1, 20) or (8, 4.5), and you plug those coordinates in for \(\Large x_1\) and \(\Large y_1\) Hopefully that makes sense?

OpenStudy (abbycross167):

oh okay

OpenStudy (abbycross167):

So like this? y-4.5=-2.214(x-8)

jimthompson5910 (jim_thompson5910):

very good, or `y - 20 = -2.214(x-1)` will work as well.

OpenStudy (abbycross167):

okay, then what?

jimthompson5910 (jim_thompson5910):

it asks to `Find the point-slope form of the equation` and you've done that when you got to `y-4.5=-2.214(x-8)` so you're done at this point

OpenStudy (abbycross167):

really? thats all?

jimthompson5910 (jim_thompson5910):

yep, what you got is in point slope form

jimthompson5910 (jim_thompson5910):